OpenStudy (anonymous):

Compare and Contrast: Below are two equations. Solve each equation and compare the two solutions. Choose the statement that is true about each solution. Equation #1: -5|3x - 2| = -15 Equation #2: |2x - 9| = 11 Equation #1 and Equation #2 have the same number of solutions. Equation #1 has more solutions than equation #2. Equation #1 has fewer solutions than equation #2. None of the statements above describe the number of solutions to the equations shown.

OpenStudy (anonymous):

i think the answer is the forst one. am i correct? @Hero @mathstudent55 @mathmath333 @jim_thompson5910

OpenStudy (anonymous):

^first one

jimthompson5910 (jim_thompson5910):

how many solutions does each equation have?

OpenStudy (anonymous):

2

jimthompson5910 (jim_thompson5910):

very good

jimthompson5910 (jim_thompson5910):

a graph can help you see this (which I'm sure you probably used)

OpenStudy (mathmath333):

\(\large\tt \begin{align} \color{black}{-5|3x - 2| = -15\\~\\ |3x - 2| = 3\\~\\ \pm(3x - 2) = 3\\~\\ x=\dfrac{5}{3}~~or~~x=-\dfrac{1}{3}\\~\\~\\ |2x - 9| = 11\\~\\ \pm(2x - 9) = 11\\~\\ x=10~~or~~x=-1}\end{align}\)

OpenStudy (anonymous):

That is exactly what I got! There are only 2 solutions for each equation right? That was all I needed to know.

jimthompson5910 (jim_thompson5910):

yes and mathmath333 shows the actual solutions and how to get them so A is correct

OpenStudy (anonymous):

Thanks to all of you!

jimthompson5910 (jim_thompson5910):

no problem